Hi, Onlie Tutors!
Here's my question.
Janie and Norbert want to sod a portion of their backyard and build the largest square concrete patio that they can afford. The area to be covered is 70 feet by 70 feet. They will do all the work themselves. The cos of sod is $0.20 per square foot. The cost of concrete is $2.00 per square foot. They have a budget of $1700. How large can the patio be?
Thank you to anyone who can help!
Amelia Brown

Answer by stanbon(75887)

Janie and Norbert want to sod a portion of their backyard and build the largest square concrete patio that they can afford. The area to be covered is 70 feet by 70 feet. They will do all the work themselves. The cost of sod is $0.20 per square foot. The cost of concrete is $2.00 per square foot. They have a budget of $1700. How large can the patio be?
------
Let the square patio be x by x ft.
Area of the patio is x^2
Cost of the patio is $2x^2
----
Area to be sodded is 70^2-x^2
Cost of the sod is 0.2(70^2-x^2)
-------
Equation:
70^2-x^2 + 0.2(70^2-x^2) <= 1700
1.2(70^2-x^2) <= 1700
70^2-x^2 <= 1416.66
x^2 >= 4900-1416.66
x^2>= 3483.33
x>= 59.02
----
Ans: The patio could be 59 ft. by 59 ft
--------
Cheers,
Stan H.
----------